OpenStudy (anonymous):

-3x^2/x+1 times x+1/6x

OpenStudy (anonymous):

\[-3x^2/(x+1) *(x+1)/(6x)=\] \[-3x^2/(6x)= -x/2\] or \[-1/2*x\]

OpenStudy (anonymous):

thanxxx

OpenStudy (anonymous):

np